# Federated Behavioural Planes: Explaining the Evolution of Client Behaviour in Federated Learning

### v1

@@ A.6 Robust aggregation baselines
A.7 Federated attacks
This section outlines and provides additional details about the federated attack scenarios implemented into our experiments, focusing on both model and data poisoning without prior knowledge of honest client data or server aggregation methods. We implemented these attacks assuming 20% of the clients are malicious:
  • Label-flipping (Data Poisoning): In a multiclass scenario, this attack changes all samples in the dataset with a source class $c_{src}$ into a target class $c_{target}$ [26, 41]. In our implementation, which involves binary classification, label flipping is performed as $1-c_{source}$ for all samples.
@@ A.7 Federated attacks
  • Crafted-noise (Model Poisoning): Inspired by the Free-riding attack described in [42], this attack aims for stealthiness by adding noise to the previous model received from the server, $w^{t}$. The noise is added as follows: $w^{*}=w^{t}+\mathcal{N}(0,\beta\cdot\sigma(w^{t}))$, where $\sigma(w^{t})$ represents the standard deviation of $w^{t}$ and $\beta$ is a scale factor. We tested different $\beta$ values: 0.3, 0.5, 0.8, 1.2, and 1.6.
@@ Appendix B Additional experiments and analysis
Client-specific adaptation in IID and non-IID scenario
As depicted in Figure 2, there is a significant reduction in the relative proximity measure between global and client-specific models across all three datasets on average, indicating a high degree of customization in the client models. Nonetheless, a few clients are still unable to tailor the model effectively to their specific distribution. To explore this phenomenon, we visually represent the generated counterfactuals pre-adaptation in Figure 5.a, showcasing models with the highest and lowest degrees of client-specific adaptation in Figures 5.c and 5.d, respectively. Notably, the most adapted model corresponds to client 4, whose data is almost perpendicular to the class boundary. In contrast, client 8, whose data distribution lies close to the boundary, struggles to adapt, remaining akin to the pre-adaptation conditions. This could highlight an increased capacity in personalisation in clients whose data is perpendicular to the learnt decision boundary, also indicating that personalization varies among clients, thereby offering more insights on the behaviour of the clients.
@@
Architecture Independence in Counterfactual Generation
To explore the independence of Federated Behavioural Planes (FBPs) from the counterfactual generator, we evaluate our methodology with an alternative model named VCNet [34], maintaining identical experimental conditions. As illustrated in Table 2, VCNet achieves comparable performance in both Federated Learning (FL) and Centralised Learning (CL) configurations, where data from all clients is placed on a single machine. Similarly to our counterfactual generator, the results demonstrate that the privacy-preserving training strategy (i.e., FL) does not compromise the VCNet’s effectiveness compared to the ideal CL condition. Operating under similar privacy constraints, FL enables VCNet to learn from the diverse distributions of client data, thereby surpassing the performance of the Local CL approach, where each client independently trains their own model.
@@
Explaining FL training with Behavioural Planes
@@
As previously observed in Figure 5, EBP and CBP provide orthogonal descriptions of client behaviors. In the Synthetic dataset, CBP effectively identifies clusters of clients with analogous data distributions. In contrast, EBP shows all honest clients moving in the same direction, thereby obscuring individual cluster distinctions but effectively contrasting the movements of honest clients with that of the attacker. Similarly, in the Breast Cancer dataset, the identification of an attack (i.e., Crafted-noise), is more discernible through EBP, which elucidates the direction of the malicious client relative to the server. These findings highlight the distinct yet complementary roles that EBP and CBP play in the analysis of client behavior across different datasets.
@@
To illustrate the complexity of model behaviour and the utility of EBP and CBP, consider the behavior of three different models with respect to the same input changes. For instance, a data point with features [1, 2] is misclassified as class 1 by Model A and correctly classified as class 0 by Model B. However, both models would change their predictions to class 1 if the features were altered to [2, 3]. In a similar fashion to Model A, a third model, Model C, may also initially misclassify the point as class 1 but would require a change in features to [2, -1] to alter its prediction to class 0, highlighting different sensitivities and prediction dynamics in response to input variations.
Federated Behavioural Shields on IID and non-IID scenarios
@@
In our main study, we focus on the non-IID setting, which is the most prevalent scenario for cross-silo FL in practice [39]. This setting presents considerable challenges due to the substantial differences between updates from honest clients, which can obscure the divergent behaviours of malicious clients [40]. Consequently, we conducted experiments on the Breast Cancer dataset in both IID and non-IID settings using our FBPs for comparative analysis. As depicted in Figure 9, we initially visualized client trajectories on both behavioural planes and within the counterfactual distance space $l$ (defined in Equation 6). The trajectories illustrate the uniform nature of honest client behaviors on both planes, distinctly highlighting the divergent behavior of the malicious client. Notably, the visualization of the counterfactual distance space, employed by our Federated Behavioural Shields, effectively identifies the malicious client. It indicates a high distance of the attacker from all other clients, leading to a low score during the aggregation process. To further validate our approach, we compared the performance of our Federated Behavioural Shields under No-attack, Crafted-noise, Inverted-gradient, Label-flipping, and Inverted-loss attacks in both IID and non-IID settings. Table 3 demonstrates that higher accuracy is achieved in the IID setting under almost all conditions compared to the non-IID setting, underscoring the increased complexity of operating in non-IID environments.
@@
Table 3: Comparison of our Federated Behavioural Shields across across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on Breast Cancer dataset for both IID and non-IID configurations
@@
Impact of server validation set size
Considering the crucial role of the validation set size on the server in the creation of behavioural planes, we initially examined its impact on computational time and accuracy through 10 distinct trials. Specifically, we evaluated our Federated Behavioural shields (without moving average) against a malicious client executing an inverted gradient attack on the Diabetes dataset. As depicted in Figure 10, computational time increases exponentially with the size of the dataset. However, a dataset containing as few as 250 samples is sufficient to adequately represent the data distribution, achieving performance on par with that observed in larger datasets. Additionally, with fewer than 1000 samples, our method demonstrates greater efficiency compared to the Krum algorithm.
@@
Ablation study
@@
Notably, relying solely on predictive performance does not yield a statistically significant benefit compared to the strongest baselines, such as Trimmed-mean on Breast Cancer and Krum on Diabetes. In contrast, counterfactual information provides deep insights into client behaviour, enabling our method to outperform all baselines in identifying malicious clients. Although the improvement is subtle, the combination of descriptors, on average, provides better results than using counterfactuals alone.
The integration of a moving average notably enhances the performance of our method, improving from $94.5\pm 0.5$ to $95.4\pm 0.3$ on the Breast Cancer dataset. Smaller gains are observed on the Diabetes dataset, likely due to two factors: the limited representativeness of the clean validation set on the server for Breast Cancer, which consists of only 89 samples compared to 500 for Diabetes, and the small size of client datasets (77 training samples on average), which may lead to unstable local training. By aggregating behaviour across multiple rounds, the moving average technique helps to stabilise and accurately assess client behaviour scores.
@@
Impact of the window length in the moving average
We analyze the impact of window length on our Federated Behavioural Shields and its two variations, one using only predictive performance (error) and the other focusing solely on decision-making processes (counterfactual). The results presented in Figure 11 depict the average accuracy across various conditions—including No attack, Label-flipping, Inverted-loss, Crafted-noise, and Inverted-gradient—for window lengths ranging from 3 to 30 rounds. While no definitive patterns are clearly evident, on average, a longer window length tends to be advantageous for accurately assessing the behavioural scores of each client during the aggregation process. Noteworthy, the optimal accuracy across both datasets was achieved with a window length of 30 rounds. Particularly, a consistent and slight increase in accuracy is observed in the Diabetes dataset when utilising only the counterfactual information.
Figure 11: Effect of window length on accuracy using Federated Behavioural Shield. This table demonstrates the accuracy changes as a function of window length in the moving average method across the Diabetes and Breast Cancer datasets.
Impact of the local epochs in Federated Behavioural Shield
Local training epochs play a crucial role in FL, particularly in evaluating client behaviours. An increased number of local epochs implies a greater adaptation of the client-model to its local data distribution. For this reason, we assess the performance of our Federated Behavioural Shields against data poisoning attacks, including Label-flipping and Inverted-loss, across Breast Cancer, Diabetes, and Synthetic datasets. Metrics are reported as the mean and standard error across ten with distinct parameters’ initialization. These attack types were selected because malicious updates are directly influenced by the number of local epochs, similar to updates from honest clients; differently, the behaviour of malicious clients in model poisoning might remain unaffected. As depicted in Figure 12, the impact of local epochs varies across datasets, with each achieving maximum accuracy at different numbers of epochs. Generally, accuracy increases with the number of local epochs until it reaches a peak, after which it begins to decline. This trend may be attributed to the fact that initially, increasing the number of local epochs enhances the detectability of anomalies in the behaviour of malicious clients, as their models diverge more quickly from those of honest clients. However, beyond a certain threshold, all client models become highly customd, leading to the potential underweighting of even honest clients’ contributions during the aggregation process on the server. This results in the system relying predominantly on a few clients whose models are the most similar to each other.
Figure 12: Impact of varying local epochs on accuracy for our Federated Behavioural Shield. This table illustrates how changes in the number of local training epochs affect accuracy across the Breast Cancer, Diabetes, and Synthetic datasets.
Federated Behavioural Shields
In this section, we provide a detailed numerical breakdown of the results depicted in Figure 4. This quantitative analysis aims to supplement the visual data presented, offering precise values and statistical insights that underpin the observations and conclusions discussed throughout the paper. Tables 5 for Breast Cancer, 6 for Diabetes, and 7 for small-MNIST present a comprehensive comparison of our Federated Behavioural Shields and its streamlined versions using only the CBP or the EBP. These are evaluated against traditional defenses such as Krum, Median, and Trimmed-mean across five scenarios: No attack, Crafted-noise, Inverted-gradient, Label-flipping, and Inverted-loss. This comparison elucidates the effectiveness of our approach under a range of conditions, both adversarial and non-adversarial.
Table 5: Comparison of Federated Behavioural Shields and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, and Trimmed-mean defenses across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on Breast Cancer dataset.
@@
Table 6: Comparison of Federated Behavioural Shields and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, and Trimmed-mean defenses across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on Diabetes dataset.
@@
Ours (error) 71.41$\pm$0.6 72.0$\pm$0.5 72.2$\pm$0.4 72.3$\pm$0.2 69.3$\pm$0.8 71.4$\pm$0.2
Ours (cf) 72.67$\pm$0.2 72.3$\pm$0.2 72.4$\pm$0.0 71.8$\pm$0.2 71.6$\pm$0.2 72.2$\pm$0.1
Ours 72.47$\pm$0.3 73.1$\pm$0.4 72.3$\pm$0.0 72.0$\pm$0.4 71.9$\pm$0.3 72.4$\pm$0.1
@@
Table 7: Comparison of Federated Behavioural Shields and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, and Trimmed-mean defenses across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on small-MNIST dataset.

### v2

@@ A.7 Robust aggregation baselines
  • Robust Federated Aggregation (RFA): RFA [42] replaces the standard arithmetic mean aggregation with the geometric median to ensure robustness against poisoned updates. It employs a Weiszfeld-type algorithm to compute the geometric median in a privacy-preserving and communication-efficient manner.
A.8 Federated attacks
This section outlines and provides additional details about the federated attack scenarios implemented into our experiments, focusing on both model and data poisoning without prior knowledge of the server aggregation methods. We implemented these attacks assuming 20% of the clients are malicious:
  • Label-flipping (Data Poisoning): In a multiclass scenario, this attack changes all samples in the dataset with a source class $c_{src}$ into a target class $c_{target}$ [26, 43]. For binary classification, label flipping is implemented as $1-c_{source}$, effectively inverting the class labels. For multiclass classification, we perform targeted label flipping by changing all instances of class 2 (Bird) to class 8 (Ship).
@@ A.8 Federated attacks
  • Crafted-noise (Model Poisoning): Inspired by the Free-riding attack described in [44], this attack aims for stealthiness by adding noise to the previous model received from the server, $w^{t}$. The noise is added as follows: $w^{*}=w^{t}+\mathcal{N}(0,\beta\cdot\sigma(w^{t}))$, where $\sigma(w^{t})$ represents the standard deviation of $w^{t}$ and $\beta$ is a scale factor. We adopted $\beta$ equal to 1.2, except when we tested the sensitivity values: 0.3, 0.5, 0.8, 1.2, and 1.6.
@@ Appendix B Additional experiments and analysis
B.1 Client-specific adaptation in IID and non-IID scenario
As depicted in Figure 2, there is a significant reduction in the relative proximity measure between global and client-specific models across all three datasets on average, indicating a high degree of customization in the client models. Nonetheless, a few clients are still unable to tailor the model effectively to their specific distribution. To explore this phenomenon, we visually represent the generated counterfactuals pre-adaptation in Figure 6.a, showcasing models with the highest and lowest degrees of client-specific adaptation in Figures 6.c and 6.d, respectively. Notably, the most adapted model corresponds to client 4, whose data is almost perpendicular to the class boundary. In contrast, client 8, whose data distribution lies close to the boundary, struggles to adapt, remaining akin to the pre-adaptation conditions. This could highlight an increased capacity in personalisation in clients whose data is perpendicular to the learnt decision boundary, also indicating that personalization varies among clients, thereby offering more insights on the behaviour of the clients.
@@ B.1 Client-specific adaptation in IID and non-IID scenario
B.2 Architecture independence in counterfactual generation
To explore the independence of FBPs from the counterfactual generator, we evaluate our methodology with an alternative model named VCNet [34], maintaining identical experimental conditions. As illustrated in Table 3, VCNet achieves comparable performance in both Federated Learning (FL) and Centralised Learning (CL) configurations, where data from all clients is placed on a single machine. Similarly to our counterfactual generator, the results demonstrate that the privacy-preserving training strategy (i.e., FL) does not compromise the VCNet’s effectiveness compared to the ideal CL condition. Operating under similar privacy constraints, FL enables VCNet to learn from the diverse distributions of client data, thereby surpassing the performance of the Local CL approach, where each client independently trains their own model.
@@ B.2 Architecture independence in counterfactual generation
B.3 Explaining FL training with Behavioural Planes
@@ B.4 The complementary roles of CBP and EBP
As previously observed in Figure 6, EBP and CBP provide orthogonal descriptions of client behaviours. In the Synthetic dataset, CBP effectively identifies clusters of clients with analogous data distributions. In contrast, EBP shows all honest clients moving in the same direction, thereby obscuring individual cluster distinctions but effectively contrasting the movements of honest clients with that of the attacker. Similarly, in the Breast Cancer dataset, the identification of an attack (i.e., Crafted-noise), is more discernible through EBP, which elucidates the direction of the malicious client relative to the server. These findings highlight the distinct yet complementary roles that EBP and CBP play in the analysis of client behaviour across different datasets.
@@
To illustrate the complexity of model behaviour and the utility of EBP and CBP, consider the behaviour of three different models with respect to the same input changes. For instance, a data point with features [1, 2] is misclassified as class 1 by Model A and correctly classified as class 0 by Model B. However, both models would change their predictions to class 1 if the features were altered to [2, 3]. In a similar fashion to Model A, a third model, Model C, may also initially misclassify the point as class 1 but would require a change in features to [2, -1] to alter its prediction to class 0, highlighting different sensitivities and prediction dynamics in response to input variations.
B.5 Client score visualization
FBPs provide a detailed representation of each client’s behaviour, enabling a deeper understanding of the conditions under which malicious clients deviate from the expected behaviour of benign clients. However, in large-scale FL systems, visual inspection may not always be necessary. Instead, statistical metrics derived from these planes can be automatically extracted and analyzed. This is demonstrated in our FBSs, which extract client-specific scores to mitigate or exclude the contributions of malicious clients during aggregation. To demonstrate the scalability and efficacy of automatic detection, we evaluated our approach using the small-CIFAR-10 dataset and recorded client scores throughout the training process using a 5-fold cross-validation scheme. As shown in Figure B.5, the extracted scores consistently identify and diminish the influence of malicious clients on the global model, as indicated by the 95% confidence interval, which varies based on the type of attack. Notably, as the model converges (with updates between rounds close to zero), more weight is given to the attacker with an inverted gradient, since its model closely resembles the global model from the previous round (plus a negligible inverted update).
((a))
((b))
((c))
((d))
Figure 10: Mean and 95% confidence interval of client scores assigned by our FBSs over 200 rounds on CIFAR-10 across different attacks.
B.6 Client scoring under unfair server-side validation sets
Our proposed methods rely on a clean validation set on the server to characterize client behaviours. However, in real-world FL scenarios, obtaining a validation set that is entirely fair and representative for each client may not be feasible. This challenge underscores the rationale behind incorporating a multi-plane evaluation in our approach, which assesses client behaviour through two distinct criteria: task performance and counterfactual analysis. Traditional performance metrics in machine learning, such as accuracy, loss, or error, are heavily dependent on the validation set used (i.e., error plane). In contrast, information derived from counterfactuals is more indicative of the learned decision process and provides insight into the training data distribution of each client, even when the validation set is biased (i.e., counterfactual plane). Consequently, an attacker or anomalous client will exhibit distinct behaviours across both planes, particularly showing an unrelated counterfactual distribution compared to other clients. Conversely, an underrepresented client will produce plausible counterfactuals similar to those of other clients, and will only be affected by the unfair validation set in the error plane.
To validate the robustness of our defense mechanism under an unfair validation set, we conducted experiments in which data from one client was excluded from the validation set. We then analyzed the behavioural scores assigned by our method. Preliminary results, shown in Figure 11, indicate that the 95% confidence interval of the scores for the underrepresented client consistently overlaps with those of other honest clients, thereby distinguishing them clearly from malicious clients.
Figure 11: Mean and 95% confidence interval of client scores assigned by our FBSs over 200 rounds on CIFAR-10 with one attacker (MP Noise) and one unfair client (i.e., their data distribution is not in the validation set).
B.7 Federated Behavioural Shields on IID and non-IID scenarios
@@
In our main study, we focus on the non-IID setting, which is the most prevalent scenario for cross-silo FL in practice [40]. This setting presents considerable challenges due to the substantial differences between updates from honest clients, which can obscure the divergent behaviours of malicious clients [41]. Consequently, we conducted experiments on the Breast Cancer dataset in both IID and non-IID settings using our FBPs for comparative analysis. As depicted in Figure 12, we initially visualised client trajectories on both behavioural planes and within the counterfactual distance space $l$ (defined in Equation 6). The trajectories illustrate the uniform nature of honest client behaviours on both planes, distinctly highlighting the divergent behaviour of the malicious client. Notably, the visualization of the counterfactual distance space, employed by our FBSs, effectively identifies the malicious client. It indicates a high distance of the attacker from all other clients, leading to a low score during the aggregation process. To further validate our approach, we compared the performance of our FBSs under No-attack, Crafted-noise, Inverted-gradient, Label-flipping, and Inverted-loss attacks in both IID and non-IID settings. Table 4 demonstrates that higher accuracy is achieved in the IID setting under almost all conditions compared to the non-IID setting, underscoring the increased complexity of operating in non-IID environments.
@@
Table 4: Comparison of our FBSs across across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on Breast Cancer dataset for both IID and non-IID configurations
@@
B.8 Impact of server validation set size
Considering the crucial role of the validation set size on the server in the creation of behavioural planes, we initially examined its impact on computational time and accuracy through 10 distinct trials. Specifically, we evaluated our FBSs (without moving average) against a malicious client executing an inverted gradient attack on the Diabetes dataset. As depicted in Figure 13, computational time increases exponentially with the size of the dataset. However, a dataset containing as few as 250 samples is sufficient to adequately represent the data distribution, achieving performance on par with that observed in larger datasets. Additionally, with fewer than 1000 samples, our method demonstrates greater efficiency compared to the Krum algorithm.
@@ B.8 Impact of server validation set size
B.9 Impact of the window length in the moving average
We analyse the impact of window length on our FBSs and its two variations, one using only predictive performance (error) and the other focusing solely on decision-making processes (counterfactual). The results presented in Figure 14 depict the average accuracy across various conditions—including No attack, Label-flipping, Inverted-loss, Crafted-noise, and Inverted-gradient—for window lengths ranging from 3 to 30 rounds. While no definitive patterns are clearly evident, on average, a longer window length tends to be advantageous for accurately assessing the behavioural scores of each client during the aggregation process. Noteworthy, the optimal accuracy across both datasets was achieved with a window length of 30 rounds. Particularly, a consistent and slight increase in accuracy is observed in the Diabetes dataset when utilising only the counterfactual information.
Figure 14: Effect of window length on accuracy using FBSs. This table demonstrates the accuracy changes as a function of window length in the moving average method across the Diabetes and Breast Cancer datasets.
B.10 Impact of the local epochs in Federated Behavioural Shields
Local training epochs play a crucial role in FL, particularly in evaluating client behaviours. An increased number of local epochs implies a greater adaptation of the client-model to its local data distribution. For this reason, we assess the performance of our FBSs against data poisoning attacks, including Label-flipping and Inverted-loss, across Breast Cancer, Diabetes, and Synthetic datasets. Metrics are reported as the mean and standard error across ten with distinct parameters’ initialization. These attack types were selected because malicious updates are directly influenced by the number of local epochs, similar to updates from honest clients; differently, the behaviour of malicious clients in model poisoning might remain unaffected. As depicted in Figure 15, the impact of local epochs varies across datasets, with each achieving maximum accuracy at different numbers of epochs. Generally, accuracy increases with the number of local epochs until it reaches a peak, after which it begins to decline. This trend may be attributed to the fact that initially, increasing the number of local epochs enhances the detectability of anomalies in the behaviour of malicious clients, as their models diverge more quickly from those of honest clients. However, beyond a certain threshold, all client models become highly customd, leading to the potential underweighting of even honest clients’ contributions during the aggregation process on the server. This results in the system relying predominantly on a few clients whose models are the most similar to each other.
Figure 15: Impact of varying local epochs on accuracy for our FBSs. This table illustrates how changes in the number of local training epochs affect accuracy across the Breast Cancer, Diabetes, and Synthetic datasets.
B.11 Impact of Attack Intensity on Federated Behavioural Shields
We conducted experiments to evaluate the sensitivity of the proposed method to varying levels of attack intensity. Specifically, we focused on model poisoning attacks by systematically increasing the noise parameter ($\beta$) injected by adversaries into the global model prior to transmission to the server ($noise=\mathcal{N}(0,\beta\cdot\sigma(w^{t}))$). The evaluation was performed using 5-fold cross-validation on the Breast Cancer and small-MNIST datasets. As shown in Figure 16, increasing the attack intensity results in a noticeable decrease in the performance of the global model when using the standard FedAvg approach. In contrast, our method—both when utilizing all feature planes and when restricted to the counterfactual plane—remains stable and unaffected by the attack intensity. Interestingly, a marginal improvement in accuracy is observed as the attack intensity increases, likely because more aggressive modifications render malicious models more degraded and, therefore, easier to detect.
((a))
((b))
Figure 16: Impact of attack intensity on FBS and traditional FedAvg in (a) the Breast Cancer dataset and (b) the small-MNIST dataset. Notably, unlike FedAvg, FBS maintains stable accuracy as the attack intensity increases.
B.12 Ablation study
@@
Notably, relying solely on predictive performance does not yield a statistically significant benefit compared to the strongest baselines, such as RFA on Breast Cancer and Krum on Diabetes. In contrast, counterfactual information provides deep insights into client behaviour, enabling our method to outperform all baselines in identifying malicious clients. Although the improvement is subtle, the combination of descriptors, on average, provides better results than using counterfactuals alone.
The integration of a moving average notably enhances the performance of our method, improving from $94.5\pm 0.5$ to $95.4\pm 0.3$ on the Breast Cancer dataset. Smaller gains are observed on the Diabetes dataset, likely due to two factors: the limited representativeness of the clean validation set on the server for Breast Cancer, which consists of only 89 samples compared to 250 for Diabetes, and the small size of client datasets (77 training samples on average), which may lead to unstable local training. By aggregating behaviour across multiple rounds, the moving average technique helps to stabilise and accurately assess client behaviour scores.
@@
B.13 Comprehensive tabular analysis of defense mechanisms
In this section, we provide a detailed numerical breakdown of the results depicted in Figure 4. This quantitative analysis aims to supplement the visual data presented, offering precise values and statistical insights that underpin the observations and conclusions discussed throughout the paper. Tables 6 for Breast Cancer, 7 for Diabetes, 8 for small-MNIST, and 9 for small-CIFAR-10 present a comprehensive comparison of our FBSs and its streamlined versions using only the CBP or the EBP. These are evaluated against traditional defenses such as Krum, Median, Trimmed-mean, RFA across five scenarios: No attack, Crafted-noise, Inverted-gradient, Label-flipping, and Inverted-loss. This comparison elucidates the effectiveness of our approach under a range of conditions, both adversarial and non-adversarial.
Table 6: Comparison of FBSs and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, and Trimmed-mean defenses across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on Breast Cancer dataset.
@@ B.13 Comprehensive tabular analysis of defense mechanisms
Table 7: Comparison of FBSs and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, and Trimmed-mean defenses across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on Diabetes dataset.
@@
RFA 70.5$\pm$0.2 69.7$\pm$0.2 69.8$\pm$0.2 65.1$\pm$0.2 62.2$\pm$0.1 67.5$\pm$0.1
Ours (error) 71.4$\pm$0.6 72.0$\pm$0.5 72.2$\pm$0.4 72.3$\pm$0.2 69.3$\pm$0.8 71.4$\pm$0.2
Ours (cf) 72.7$\pm$0.2 72.3$\pm$0.2 72.4$\pm$0.0 71.8$\pm$0.2 71.6$\pm$0.2 72.2$\pm$0.1
Ours 72.5$\pm$0.3 73.1$\pm$0.4 72.3$\pm$0.0 72.0$\pm$0.4 71.9$\pm$0.3 72.4$\pm$0.1
@@
Table 8: Comparison of FBSs and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, and Trimmed-mean defenses across five attacks types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on small-MNIST dataset.
@@
Table 9: Comparison of FBSs and its simpler version with only counterfactuals (cf) or predictive-performance (error) versus Krum, Median, Trimmed-mean, and RFA defenses across five attack types—No attack, Crafted-noise, Inverted-gradient, Label-flipping, Inverted-loss—on small-CIFAR-10 dataset.
Model Poisoning Data Poisoning
No-Attack Crafted-Noise Inv. Grad. Label-Flip Inv. Loss Mean
Krum 90.6$\pm$2.0 90.7$\pm$1.3 92.1$\pm$1.1 93.0$\pm$1.3 85.3$\pm$2.4 90.3$\pm$0.3
Median 96.5$\pm$0.4 97.0$\pm$0.4 96.1$\pm$0.2 96.0$\pm$0.7 80.6$\pm$5.3 93.2$\pm$0.5
Trim 96.5$\pm$0.3 96.9$\pm$0.3 97.4$\pm$0.3 96.2$\pm$0.6 86.3$\pm$3.4 94.7$\pm$0.3
RFA 96.3$\pm$0.4 97.2$\pm$0.3 97.0$\pm$0.3 96.3$\pm$0.8 83.8$\pm$3.6 94.1$\pm$0.3
Ours (error) 97.4$\pm$0.2 97.1$\pm$0.2 96.1$\pm$0.4 97.3$\pm$0.5 89.8$\pm$1.2 95.5$\pm$0.1
Ours (cf) 98.2$\pm$0.2 97.4$\pm$0.3 97.6$\pm$0.1 97.6$\pm$0.2 91.8$\pm$0.8 96.5$\pm$0.1
Ours 97.4$\pm$0.3 97.9$\pm$0.4 97.3$\pm$0.3 97.2$\pm$0.4 91.7$\pm$1.0 96.3$\pm$0.1
Predictor 97.5$\pm$0.3 N/A N/A N/A N/A N/A